OpenStudy (anonymous):

Graph f(x) = |3x+2| + |2-3x|

OpenStudy (anonymous):

I3x+2I = 3x+2 , x>-2/3 = - (3x+2) , x<-2/3 I2-3xI = 2-3x , x< 2/3 = -(2-3x) , x>2/3 you can now write the fn in three ranges 1) from negative infinity to -2/3 2) -2/3 to 2/3 3) 2/3 to infinty

OpenStudy (anonymous):

so the following equation should be = 3x+2 - (2-3x) for x >= 1, 3x+2+2-3x for -2/3 <= x <= 2/3, -(3x+2) + 2 - 3x for x < -2/3

OpenStudy (anonymous):

which equals to 6x for x>= 2/3 4 fpr -2/3 <= x < 2/3 -6x for x< - 2/3

OpenStudy (anonymous):